# A Variant of the Bravyi-Terhal Bound for Arbitrary Boundary Conditions

### v1

@@ 4 Minimum distance bound for Abelian 2BGA codes
Proof: Let $\mathcal{C}$ be a non-trivial Abelian 2BGA code defined by $\mathbf{A}=\sum_{g\in G}a_{g}\mathbb{B}(g)$ and $\mathbf{B}=\sum_{g\in G}b_{g}\mathbb{B}(g)$ associated to the non-zero elements $a=\sum_{g\in G}a_{g}g$ and $b=\sum_{g\in G}b_{g}g$ of the algebra group $\mathbb{F}_{2}[G]$. Consider $g_{a},g_{b}\in G$ s.t. $a_{g_{a}}=b_{g_{b}}=1$. Consider $\tilde{a}=g_{a}^{-1}a=e+\sum_{g\in G\setminus\{g_{a}\}}a_{g}g$ and $\tilde{b}=g_{b}^{-1}b=e+\sum_{g\in G\setminus\{g_{b}\}}b_{g}g$. The matrices constructed from $\tilde{a}$ and $\tilde{b}$ are respectively the matrices $\tilde{\mathbf{A}}=\sum_{g\in G}a_{g}\mathbb{B}(g_{a}^{-1}g)=\mathbb{B}(g_{a}^{-1})\sum_{g\in G}a_{g}\mathbb{B}(g)=\mathbb{B}(g_{a}^{-1})\mathbf{A}$ and $\tilde{\mathbf{B}}=\mathbb{B}(g_{b}^{-1})\mathbf{B}$. The endomorphism of $\mathbb{F}_{2}^{2n}$ given by $(u\hskip 5.69046ptv)\mapsto(\mathbb{B}(g_{a}^{-1})u\hskip 5.69046pt\mathbb{B}(g_{b}^{-1})v)$ is an ismorphism that preserves the Hamming metric and induces bijections between the kernels of $[\mathbf{A},\,\mathbf{B}]$ and $[\mathbb{B}(g_{a}^{-1})\mathbf{A},\,\mathbb{B}(g_{b}^{-1})\mathbf{B}]$, and between the row-spaces of $[\mathbf{B}^{\intercal},\,\mathbf{A}^{\intercal}]$ and $[(\mathbb{B}(g_{b}^{-1})\mathbf{B})^{\intercal},\,(\mathbb{B}(g_{a}^{-1})\mathbf{A})^{\intercal}]$. Hence the Abelian 2BGA codes defined by $(\mathbf{A},\mathbf{B})$ and $(\tilde{\mathbf{A}},\tilde{\mathbf{B}})$ are equivalent. Consequently, to study the minimum distance of the code, we may assume that $e\in\mathrm{supp}(a)$ and $e\in\mathrm{supp}(b)$. . $\square$
We may now suppose that $\mathrm{supp}(a)=\{g\in G\,|\,a_{g}\neq 0\}=\{e,g_{a_{1}},\ldots,g_{a_{r}}\}$ and $\mathrm{supp}(b)=\{e,g_{b_{1}},\ldots,g_{b_{s}}\}$. Consider now the $\mathbb{Z}$-linear map
$\Psi:\mathbb{Z}^{r}\times\mathbb{Z}^{s}\rightarrow G,\;\epsilon_{i}\mapsto\left\{\begin{array}[]{ll}g_{a_{i}}&\text{if }1\leq i\leq r,\\ g_{b_{i-r}}&\text{if }r+1\leq i\leq r+s\end{array}\right.,$
@@
Proof: Consider the subgroup given by the image of $\Psi$, $H:=\mathrm{Im}(\Psi)=\prod_{i=1}^{r}<g_{a_{i}}>\times\prod_{i=1}^{s}<g_{b_{i}}>\,\subset G$. If $H\neq G$, then the code is a copy of $[G:H]$ codes of length $2|H|$. Indeed, one can reorder the elements of $G$ in such a way that they represent the successive cosets of $G/H$. Since $a,b\in\mathbb{F}_{2}[H]$, $\forall g\in G$, $ga,gb\in\mathbb{F}_{2}[gH]$. Hence $\mathbf{A}$ and $\mathbf{B}$ are block diagonal matrices and so by considering each pair of blocks of $\mathbf{A}$ and $\mathbf{B}$ individually, we obtain $[G:H]$ equivalent 2BGA codes of length $2|H|$. . $\square$
@@
Proof: For $i\in[[0,n-1]]$, the support of the $i$-th $X$ generator is $S^{X}_{i}=\{\bar{\Psi}^{-1}(g_{j})|(\mathbf{H}_{X})_{i,j}=1\text{ or }(\mathbf{H}_{X})_{i,n+j}=1\}$ where $j\in[[0,n-1]]$. This generator is indexed by the $i$-th row of $\mathbf{A}$ and $\mathbf{B}$ and so
  • $\{g_{j}\,|j\in[[0,n-1]],(\mathbf{H}_{X})_{i,j}=1\}=\{g_{i},g_{a_{1}}^{-1}g_{i},\ldots,g_{a_{r}}^{-1}g_{i}\}$
@@
Hence, it follows that $S^{X}_{i}=\{\bar{\Psi}^{-1}(g_{i}),\bar{\Psi}^{-1}(g_{i})-\epsilon_{1},\ldots,\bar{\Psi}^{-1}(g_{i})-\epsilon_{r+s}\})$ which is included in $B(\bar{\Psi}^{-1}(g_{i}),1)$.
Similarly, the support of the $i$-th $Z$ generator is $S^{Z}_{i}=\{\bar{\Psi}^{-1}(g_{j})|(\mathbf{H}_{Z})_{i,j}=1\text{ or }(\mathbf{H}_{Z})_{i,n+j}=1\}$ which is indexed by the $i$-th column of $\mathbf{B}$ and $\mathbf{A}$ and so
  • $\{g_{j}|j\in[[0,n-1]],\,(\mathbf{H}_{Z})_{i,j}=1\}=\{g_{i},g_{b_{1}}g_{i},\ldots,g_{b_{s}}g_{i}\}$
  • $\{g_{j}|j\in[[0,n-1]],\,(\mathbf{H}_{Z})_{i,n+j}=1\}=\{g_{i},g_{a_{1}}g_{i},\ldots,g_{a_{r}}g_{i}\}$

### v2

@@ 4 Minimum distance bound for Abelian 2BGA codes
Proof: Let $\mathcal{C}$ be a non-trivial Abelian 2BGA code defined by the matrices $\mathbf{A}=\sum_{g\in G}a_{g}\mathbb{B}(g)$ and $\mathbf{B}=\sum_{g\in G}b_{g}\mathbb{B}(g)$ associated to the non-zero elements $a=\sum_{g\in G}a_{g}g$ and $b=\sum_{g\in G}b_{g}g$ of the algebra group $\mathbb{F}_{2}[G]$. Consider $g_{a},g_{b}\in G$ s.t. $a_{g_{a}}=b_{g_{b}}=1$. We now consider the elements $\tilde{a}=g_{a}^{-1}a=e+\sum_{g\in G\setminus\{g_{a}\}}a_{g}g$ and $\tilde{b}=g_{b}^{-1}b=e+\sum_{g\in G\setminus\{g_{b}\}}b_{g}g$. The matrices constructed from $\tilde{a}$ and $\tilde{b}$ are respectively the matrices $\tilde{\mathbf{A}}=\sum_{g\in G}a_{g}\mathbb{B}(g_{a}^{-1}g)=\mathbb{B}(g_{a}^{-1})\sum_{g\in G}a_{g}\mathbb{B}(g)=\mathbb{B}(g_{a}^{-1})\mathbf{A}$ and $\tilde{\mathbf{B}}=\mathbb{B}(g_{b}^{-1})\mathbf{B}$. The endomorphism of $\mathbb{F}_{2}^{2n}$ given by $(u\hskip 5.69046ptv)\mapsto(\mathbb{B}(g_{a}^{-1})u\hskip 5.69046pt\mathbb{B}(g_{b}^{-1})v)$ is an ismorphism that preserves the Hamming metric and induces bijections between the kernels of $[\mathbf{A},\,\mathbf{B}]$ and $[\mathbb{B}(g_{a}^{-1})\mathbf{A},\,\mathbb{B}(g_{b}^{-1})\mathbf{B}]$, and between the row-spaces of $[\mathbf{B}^{\intercal},\,\mathbf{A}^{\intercal}]$ and $[(\mathbb{B}(g_{b}^{-1})\mathbf{B})^{\intercal},\,(\mathbb{B}(g_{a}^{-1})\mathbf{A})^{\intercal}]$. Hence the Abelian 2BGA codes defined by $(\mathbf{A},\mathbf{B})$ and $(\tilde{\mathbf{A}},\tilde{\mathbf{B}})$ are equivalent. Consequently, to study the minimum distance of the code, we may assume that $e\in\mathrm{supp}(a)$ and $e\in\mathrm{supp}(b)$. . $\square$
We may now suppose that $\mathrm{supp}(a)=\{g\in G\,|\,a_{g}\neq 0\}=\{e,g_{a_{1}},\ldots,g_{a_{r}}\}$ and similarly that $\mathrm{supp}(b)=\{e,g_{b_{1}},\ldots,g_{b_{s}}\}$. Consider now the $\mathbb{Z}$-linear map
$\Psi:\mathbb{Z}^{r}\times\mathbb{Z}^{s}\rightarrow G,\;\epsilon_{i}\mapsto\left\{\begin{array}[]{ll}g_{a_{i}}&\text{if }1\leq i\leq r,\\ g_{b_{i-r}}&\text{if }r+1\leq i\leq r+s,\end{array}\right.$
@@
Proof: Consider the subgroup given by the image of $\Psi$, $H:=\mathrm{Im}(\Psi)\,\subset G$. If $H\neq G$, then the code is a copy of $[G:H]$ codes of length $2|H|$. Indeed, one can reorder the elements of $G$ in such a way that they represent the successive cosets of $G/H$. Since $a,b\in\mathbb{F}_{2}[H]$, $\forall g\in G$, $ga,gb\in\mathbb{F}_{2}[gH]$. Hence $\mathbf{A}$ and $\mathbf{B}$ are block diagonal matrices and so by considering each pair of blocks of $\mathbf{A}$ and $\mathbf{B}$ individually, we obtain $[G:H]$ equivalent 2BGA codes of length $2|H|$. . $\square$
@@
Proof: For $i\in[[0,n-1]]$, the support of the $i$-th $X$-generator is $S^{X}_{i}$ which is indexed by the $i$-th row of $\mathbf{A}$ and $\mathbf{B}$. From the definition of the permutation matrices $\mathbb{B}(g)$, the support $S^{X}_{i}$ is thus given as the image of $\bar{\Psi}^{-1}$ of the two following sets:
  • $\{g_{j}\,|j\in[[0,n-1]],(\mathbf{H}_{X})_{i,j}=1\}=\{g_{i},g_{a_{1}}^{-1}g_{i},\ldots,g_{a_{r}}^{-1}g_{i}\}$,
@@
Hence, it follows that $S^{X}_{i}=\{\bar{\Psi}^{-1}(g_{i}),\bar{\Psi}^{-1}(g_{i})-\epsilon_{1},\ldots,\bar{\Psi}^{-1}(g_{i})-\epsilon_{r+s}\})$ which is included in $B(\bar{\Psi}^{-1}(g_{i}),1)$. Similarly, the support of the $i$-th $Z$-generator is $S^{Z}_{i}$ which is indexed by the $i$-th column of $\mathbf{B}$ and $\mathbf{A}$. From the definition of the permutation matrices $\mathbb{B}(g)$, the support $S^{Z}_{i}$ is thus given as the image of $\bar{\Psi}^{-1}$ of the two following sets:
  • $\{g_{j}|j\in[[0,n-1]],\,(\mathbf{H}_{Z})_{i,j}=1\}=\{g_{i},g_{b_{1}}g_{i},\ldots,g_{b_{s}}g_{i}\}$,
  • $\{g_{j}|j\in[[0,n-1]],\,(\mathbf{H}_{Z})_{i,n+j}=1\}=\{g_{i},g_{a_{1}}g_{i},\ldots,g_{a_{r}}g_{i}\}$.